OpenStudy (anonymous):

The histogram below shows the number of hoagie sandwiches students in Carson's class sold for their fundraiser. Use the histogram to match each conclusion with its value. Round to the nearest whole number if necessary.

OpenStudy (anonymous):

percent of the class that sold less than twenty hoagies A) forty-nine greatest number of hoagies that could have been sold by one student B) twenty total number of students in Carson's class C) thirty-two more than half of the students sold at least this number of hoagies D) twenty-two

OpenStudy (anonymous):

@osamabinswaggin

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

i dont see the histogram

OpenStudy (anonymous):

okay here I got it

OpenStudy (anonymous):

is it hard :P

OpenStudy (anonymous):

@triciaal is it hard :P

OpenStudy (anonymous):

please do you have the answer I want to finish

OpenStudy (triciaal):

sorry i was away I will look at it now

OpenStudy (triciaal):

22 students in the class

OpenStudy (triciaal):

(2 + 5) = 7 sold less than 20 percent means per 100 7/22 = x/100 solve for x x = 32 32% of class sold less than 20

OpenStudy (triciaal):

greatest number sold by one student = 49